The sum of radii of two circles is 10.5 cm and the difference of their circumferences is 13.2 cm. Find the radii of the two circles.
Answer by TimothyLamb(4379)

c = 2pi*r
---
m = radius of smaller circle
n = radius of larger circle
---
m + n = 10.5
m = 10.5 - n
2pi*n - 2pi*m = 13.2
---
2pi*n - 2pi*m = 13.2
n - m = 13.2/2pi
n - m = 2.1008452
n - (10.5 - n) = 2.1008452
n - 10.5 + n = 2.1008452
2n = 12.600845
---
n = 6.3004226
m = 4.1995774
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php